Can I get help to solve this question:
Find all the values of & x in the int(0,2pi) & for the equation sinx = -cosx
1
Expert's answer
2011-08-15T06:22:33-0400
sinx = -cos x
tan x = -1
x = 3/4 π + π n, n = 0,1
I.e.
x1 = 3/4 π
x2 = 7/4 π